THE S_{1} ← S_{0} FLUORESCENCE EXCITATION SPECTRA OF ANISOLE AND STYRENE USING HIGH RESOLUTION TECHNIQUES

\underline{JASON W. RIBBLETT} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, PA 15260.

Anisole (methoxybenzene) and styrene (vinylbenzene) are similar chemical species. However, their electronic structures are very different. In this report, we describe the rotationally resolved fluorescence excitation spectra of the S_{1} ← S_{0} origin bands of anisole and styrene. Both spectra are similar in overall appearance. However, a detailed analysis of the two spectra shows that the S_{1} ← S_{0} optical transition moments have quite different orientations in the respective molecular frames. The results will be interpreted in the framework of a state-mixing model for the S_{1} states of the two molecules.